A cyclon fence will cover a rectangular piece of land, which area measures 800fts cubics and the height of the land double its wide. How much feets of cyclon fence will need?
Answer by edjones(8007)

A=area L=length W=width
A=L*W
L=2W
substitute 2W for L: A=W*2W
W*2W=800
2W^2=800
divide both by 2: W^2=400
square root of both: W=20 feet
L=40 feet
Check:
20*40=800 square feet
Ed